Slope of a budget line = - price of good 2 / price of good 1.

Healthy Meals and Unhealthy Snacks

Budget

Price of Healthy Meals (Good 1)

Price of Unhealthy Snacks (Good 2)

Scenario 1

$80.00

$5.00

$2.00

Scenario 2

$80.00

$5.00

$1.25

Scenario 3

$120.00

$5.00

$2.00

  1. What is the X intercept?
  2. What is the Y intercept?
  3. What is the slope?
Answer #1

Scenario 1:- X intercept = 80/5 = 16, Y intercept = 80/2 =40, Slope = 2/5 = 0.4

Scenario 2:- X intercept = 80/5 =16, Y intercept = 80/1.25 = 64, Slope = 1.25/5 = 0.25

Scenario3:- X intercept = 120/5 = 24, Y intercept = 120/2 = 60, Slope = 2/5 = 0.4